# Bayes2IMC: In-Memory Computing for Bayesian Binary Neural Networks

### v1

@@ Abstract
Bayesian Neural Networks (BNNs) provide superior estimates of uncertainty by generating an ensemble of predictive distributions. However, inference via ensembling is resource-intensive, requiring additional entropy sources to generate stochasticity which increases resource consumption. In this work, we introduce Bayes2IMC, an in-memory computing (IMC) architecture designed for binary Bayesian neural networks that leverage nanoscale device stochasticity to generate desired distributions. Our novel approach utilizes Phase-Change Memory (PCM) to harness inherent noise characteristics, enabling the creation of a binary neural network. This design eliminates the necessity for a pre-neuron Analog-to-Digital Converter (ADC), significantly improving power and area efficiency. We also develop a hardware-software co-optimized correction method applied solely on the logits in the final layer to reduce device-induced accuracy variations across deployments on hardware. Additionally, we devise a simple compensation technique that ensures no drop in classification accuracy despite conductance drift of PCM. We validate the effectiveness of our approach on the CIFAR-10 dataset with a VGGBinaryConnect model containing 14 million parameters, achieving accuracy metrics comparable to ideal software implementations as well as results reported in the literature using other technologies. Finally, we present a complete core architecture and compare its projected power, performance, and area efficiency against an equivalent SRAM baseline, showing a $3.8$ to 9.6×9.6$\times$9.6 × improvement in total efficiency (in GOPS/W/mm2) and a $2.2$ to 5.6×5.6$\times$5.6 × improvement in power efficiency (in GOPS/W). In addition, the projected hardware performance of Bayes2IMC surpasses that of most of the BNN architectures based on memristive devices reported in the literature, and achieves up to $20\%$ higher power efficiency compared to the state-of-the-art.
@@ I-A Context and Motivation
The growing demand for edge AI in applications such as medical diagnostics [1], facial identification and surveillance [2] and self-driving cars [3] where reliability and safety are of paramount importance, has heightened the need for systems that provide a measure of uncertainty while operating under significant resource constraints in terms of area and power. Advances in deep neural networks have ensured very high accuracy, but this has come at the cost of overconfidence and poor calibration [4]. More importantly, such networks are incapable of uncertainty quantification, especially out-of-distribution identification measured by epistemic uncertainty.[5]. Bayesian neural networks (BNNs) overcome this problem by allowing the weight parameters to be probability distributions rather than point estimates, thus encoding uncertainty in the parameter distributions [6, 7]. During inference, the network weights are then sampled to create an ensemble, and the outputs are combined in Monte-Carlo (MC) fashion to obtain predictions, confidence and uncertainty (Fig. 1). The creation of ensembles can be done in time, i.e., instantiating each ensemble one after another on the same hardware [8], or in space, i.e., instantiating all the ensembles simultaneously, using multiple copies of the hardware [6]. Both methods are resource-intensive compared to a traditional network whose parameters are point estimates and require only a single instantiation. The ensembling-in-time approach reduces net decision-making throughput while optimizing for power and total area. In contrast, the ensembling-in-space expends more area and power in order to deliver high net throughput. In addition, both these approaches require noise sources to generate random numbers. Thus, the deployment of Bayesian neural networks at the edge is challenging and requires careful hardware-software co-optimization [9, 10].
@@
In this paper, we design and validate an area and power-efficient ensembling-in-time Bayesian binary neural netwrok in-memory compute (IMC) architecture named Bayes2IMC that leverages the inherent stochasticity of nanoscale memory devices. Fig. 1 provides a high-level description of this core. The nanoscale devices are arranged in a crossbar fashion, which is partitioned into two sections: a weight plane (WP) that stores the distribution parameters, and a noise plane (NP) that provides stochasticity necessary for sampling [6]. We also modify the traditional in-memory computing approach by routing inputs to an accumulator for multiply-and-accumulate (MAC) operations. This combination of routing strategy, row-by-row read flow, and on-the-fly sample generation during sensing enables inference without the need for an analog-to-digital converter, thereby improving area and power efficiency.
@@ I-B Background and Related Work
In-Memory Computing: The IMC approach has been proposed to address the ‘von Neumann bottleneck’ [11] plaguing traditional computing platforms [12, 13, 14, 15, 16]. In such systems, certain computational operations are performed in memory without data movement, enabling high throughput at low area and power. IMC implementations based on digital Static Random Access Memories (SRAMs) [17, 18, 16, 19, 20] as well as those based on non-volatile memory (NVM) devices such as Phase Change Memory (PCMs), Resistive RAMs (RRAMs), Spin-Transfer Torque RAMs (STT-RAMs), and Spin-Orbit Torque Magnetic RAMs (SOT-MRAMs) have been explored for implementing frequentist deep learning models [21, 14, 22]. Compared to SRAM IMC implementations, NVM devices have following advantages:
  • Multi-level conductance for multi-bit parameter storage.
  • Near $\mathcal{O}(1)$ matrix-vector multiplication (MVM) operation when arranged in crossbar fashion with input vector encoded as drive voltage [23].
However, NVM devices also pose several challenges for IMC MVM implementation. The high level of state-dependent programming [24] results in write error, and conductance drift results in instability of programmed weights [24, 10]. This severely impacts the overall classification accuracy. Additional sources of error include MVM errors due to IR drops on source-line due to high current in the all-parallel operation mode [18], and non-linear dependence of device current on applied voltage [25]. Furthermore, Analog-to-Digital converters (ADCs) are necessary to convert the analog output of the MVM operation to multi-bit digital values, and this impacts both the overall power consumption and area efficiency [18, 12].
Related works: We now discuss proposed IMC solutions in the literature that leverage the inherent stochasticity of NVM devices with the goal of implementing ensembling methods. An RRAM array-based Bayesian network with training implemented via Markov Chain Monte Carlo (MCMC) was demonstrated in [26], though the need for training in situ is time and resource intensive [27]. Bernoulli distributed noise of Domain Wall Magnetic Tunnel Junction (DW-MTJ) devices and SOT-MRAM were used in [21] and [9] respectively to generate Gaussian random variables. The latter uses a local reparametrization technique to reduce resource consumption. However, both of these works require the mean and standard deviation parameters to be stored separately and the contributions from multiple devices to be combined to generate a single random variate per the central limit theorem. This lead to expensive hardware requirements.
@@ I-C Main Contributions
  • By introducing row-by-row read, we avoid high output currents. Additionally, every read operation handles the binarization in tandem without requiring any throughput throttling step. This permits an ADC-less architecture, significantly improving the efficiency of the design.
  • We avoid using DACs at every row by streaming inputs directly into pre-neuron accumulator and applying fixed width and amplitude read pulses on bit-lines to sample weights. By doing so, we improve the area and power efficiency and avoid the MVM errors due to the non-linear dependence of device conductance on applied voltages.
@@ II-A Binary Bayesian Neural Networks
Given a parameterized predictor $p(y|\boldsymbol{x},\boldsymbol{w})$ with parameter vector $\boldsymbol{w}\in\mathcal{W}$, Bayesian learning amounts to the evaluation of the posterior distribution $p(\boldsymbol{w}|\mathcal{D})$, quantifying the likelihood of each parameter $\boldsymbol{w}$ under the evidence presented by the training dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i}$. At inference time, the contribution of each model $p(y|\boldsymbol{x},\boldsymbol{w})$ is weighted by the posterior $p(\boldsymbol{w}|\mathcal{D})$ [32, 30], yielding the marginal distribution
[…]
In this work, we focus on the implementation of BBNNs, where each weight $w\in\boldsymbol{w}$ of the neural network $\mathbf{f}_{\boldsymbol{w}}(\boldsymbol{x})=\{p(y|\boldsymbol{x},\boldsymbol{w})\}_{y}$ is restricted to a binary set of values $w\in\{-1,1\}$. Accordingly, the parameter space is defined as $\mathcal{W}=\{-1,1\}^{|\boldsymbol{w}|}$, where $|\boldsymbol{w}|$ represents the number of synaptic weights. Intuitively, BBNNs implement a set of neural networks $\{\mathbf{f}_{\boldsymbol{w}}\}_{\boldsymbol{w}\sim p(\boldsymbol{w}|\mathcal{D})}$ where the weights are trained to follow a posterior distribution $p(\boldsymbol{w}|\mathcal{D})$ rather than a single point-estimate [30]. Bayesian learning for BBNN is performed through mean-field variational inference (VI) [33, 34]. The posterior distribution $p(\boldsymbol{w}|\mathcal{D})$ is approximated by a parameterized distribution $q(\boldsymbol{w}|\boldsymbol{\lambda})$, where $\boldsymbol{\lambda}$, the variational parameter, is the vector of natural parameters of Bernoulli distribution $\lambda_{w}$ in the exponential family, i.e $\boldsymbol{\lambda}=\{\lambda_{w}\}_{w\in\boldsymbol{w}}\in\mathbb{R}^{|\boldsymbol{w}|}$. The probability parameter $p_{w}$ describing the distribution of binary weights $w$ is encoded as
@@
After training, we obtain a collection of synaptic probabilities $\boldsymbol{p}_{\boldsymbol{w}}=\{p_{w}\}_{w\in\boldsymbol{w}}$ from (2) which, using an appropriate binary random variable generator, can be leveraged to create an ensemble of networks $\mathbf{f}_{\boldsymbol{w}}$ with $\boldsymbol{w}\sim q(\boldsymbol{w}|\boldsymbol{\lambda})$. As illustrated in Fig. 1, during inference, the outputs $p(y|\boldsymbol{x},\boldsymbol{w})$ of the sampled networks $\mathbf{\hat{f}}_{\boldsymbol{w}}$ are accumulated and processed to obtain the final marginal predictor
$p(y|\boldsymbol{x},\mathcal{D})=\mathbb{E}_{\boldsymbol{w}\sim q(\boldsymbol{w}|\boldsymbol{\lambda})}\left[p(y|\boldsymbol{x},\boldsymbol{w})\right]\approx\frac{1}{N_{MC}}\sum_{i=1}^{N_{MC}}p(y|\boldsymbol{x},\boldsymbol{w}_{i}),$ (4)
@@ II-B Expected calibration error (ECE)
The ECE quantifies the in-distribution calibration of a trained network by quantifying the expected discrepancy between accuracy and the confidence scores obtained by the output probability of the predicted class. For a given dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N_{s}}$ of $N_{s}$ samples, we denote as $B_{m}=\{(\hat{p}_{i},\hat{y}_{i},y_{i})\}_{i|\hat{p}_{i}\in I_{m}}$ the set of confidence-label pairs for which the predicted confidence $\hat{p}_{i}=p(y_{i}|x_{i},\mathcal{D})$ lies inside $I_{m}$, with $m\in\{1,...,N_{b}\}$. ECE is mathematically described, as per [4], as
@@ II-C Aleatoric and Epistemic Uncertainty
where $\{\boldsymbol{w}_{i}\}_{i=1}^{N_{MC}}\overset{\mathrm{i.i.d.}}{\sim}q(\boldsymbol{w}|\boldsymbol{\lambda})$ and $p(y|\boldsymbol{x},\boldsymbol{w}_{i})$ is the probability value for the given class $y\in\mathscr{C}$. This total uncertainty $U_{tot}$ can be decomposed into the uncertainty due to the inherently stochastic nature of the covariate-label pairs $(\boldsymbol{x},y)$, i.e., the aleatoric uncertainty; and the uncertainty stemming from the limited amount of evidence $\mathcal{D}$, i.e., the epistemic uncertainty.
@@
Ideally, for an out-of-distribution (OOD) data, the ensemble of predictors will disagree on predictions, causing a large epistemic uncertainty. In contrast, on In-distribution (IND) data, there should be a broad agreement on the prediction probabilities, and thus the epistemic uncertainty should be low. Fig. 2 illustrates the distinct roles that aleatoric and epistemic uncertainties play in understanding the predictions of a BBNN. Fig. 2 and Fig. 2 consider the output of the BBNN to IND data. for inputs characteriszed by a higher confidence in the prediction hence a lower aleatoric uncertainty. The higher aleatoric in the latter case suggests low prediction confidence. In addition, and by a lower prediction confidence hence higher aleatoric uncertainty respectively. In addition, the epistemic uncertainty is low in both cases. This suggests that the incorrect prediction is due to data uncertainty rather than the model’s lack of confidence in its predictions. Fig. 2 and Fig. 2 consider OOD data, in which case the high epistemic uncertainties indicate the disagreement amongst the ensembles due to the unfamiliar nature of the inputs.
@@ III Architecture and Design
In this section, we discuss the features of Bayes2IMC, our proposed architecture illustrated in Fig. 1. We first introduce the mathematical techniques that allow the utilization of device stochasticity as a source of noise distribution in Sec. III-A and the corresponding hardware implementation in Sec. III-B. We then describe the details the core operations of the proposed Bayes2IMC architecture (Sec. III-C) and introduce the proposed output logit correction scheme (Sec. III-D). Finally, Sec III-E discusses the drift compensation mechanism.

### v2

@@ Abstract
Bayesian Neural Networks (BNNs) generate an ensemble of possible models by treating model weights as random variables. This enables them to provide superior estimates of decision uncertainty. However, implementing Bayesian inference in hardware is resource-intensive, as it requires noise sources to generate the desired model weights. In this work, we introduce Bayes2IMC, an in-memory computing (IMC) architecture designed for binary BNNs that leverages the stochasticity inherent to nanoscale devices. Our novel design, based on Phase-Change Memory (PCM) crossbar arrays eliminates the necessity for Analog-to-Digital Converter (ADC) within the array, significantly improving power and area efficiency. Hardware-software co-optimized corrections are introduced to reduce device-induced accuracy variations across deployments on hardware, as well as to mitigate the effect of conductance drift of PCM devices. We validate the effectiveness of our approach on the CIFAR-10 dataset with a VGGBinaryConnect model containing 14 million parameters, achieving accuracy metrics comparable to ideal software implementations. We also present a complete core architecture, and compare its projected power, performance, and area efficiency against an equivalent SRAM baseline, showing a $3.8$ to 9.6×9.6$\times$9.6 × improvement in total efficiency (in GOPS/W/mm2) and a $2.2$ to 5.6×5.6$\times$5.6 × improvement in power efficiency (in GOPS/W). In addition, the projected hardware performance of Bayes2IMC surpasses most memristive BNN architectures reported in the literature, achieving up to $20\%$ higher power efficiency compared to the state-of-the-art.
@@ I-A Context and Motivation
The growing demand for edge AI in applications such as medical diagnostics [1], facial identification and surveillance [2] and self-driving cars [3] where reliability and safety are of paramount importance, has heightened the need for systems that provide a measure of uncertainty while operating under significant resource constraints in terms of area and power. Advances in deep neural networks have ensured very high accuracy, but this has come at the cost of overconfidence and poor calibration [4]. Specifically, such networks are incapable of distinguishing between aleatoric uncertainty (data uncertainty) and epistemic uncertainty (model uncertainty),[5], making tasks such as out-of-distribution (OOD) detection particularly challenging. Bayesian neural networks (BNNs) overcome this problem by allowing the weights to be random variables rather than point estimates, thus encoding uncertainty in the parameter distributions [6, 7]. To perform inference, the network weights are sampled to create an ensemble of predictors, whose outputs are then combined to obtain data class prediction and uncertainty estimates (Fig. 1).
The creation of ensembles can be done in time, i.e., instantiating each ensemble one after another on the same hardware [8], or in space, i.e., instantiating all the ensembles simultaneously, using multiple copies of the hardware [6]. Both methods are resource-intensive compared to a traditional network whose parameters are point estimates and require only a single instantiation.
The ensembling-in-time approach reduces net decision-making throughput while optimizing for power and total area. In contrast, the ensembling-in-space expends more area and power in order to deliver high net throughput. In addition, both these approaches require noise sources to generate random numbers. Thus, the deployment of Bayesian neural networks at the edge is challenging and requires careful hardware-software co-optimization [9, 10].
@@
In this paper, we design and validate an area and power-efficient ensembling-in-time Bayesian binary neural network in-memory compute (IMC) architecture named Bayes2IMC that leverages the inherent stochasticity of nanoscale memory devices. Fig. 1 provides a high-level description of this core. The nanoscale devices are arranged in a crossbar fashion, which is partitioned into two sections: a weight plane (WP) that stores the distribution parameters, and a noise plane (NP) that provides stochasticity necessary for sampling. This approach was first introduced in the conference version [6] of this work. We also modify the traditional in-memory computing approach by routing inputs to an accumulator for multiply-and-accumulate (MAC) operations. This combination of routing strategy, row-by-row read flow, and on-the-fly sample generation during sensing enables inference without the need for an analog-to-digital converter, thereby improving area and power efficiency.
@@ I-B Background and Related Work
In-Memory Computing: The IMC approach has been proposed to address the ‘von Neumann bottleneck’ [11] plaguing traditional computing platforms [12, 13, 14, 15, 16]. In such systems, certain computational operations are performed in memory without data movement, enabling high throughput at low area and power. IMC implementations based on digital Static Random Access Memories (SRAMs) [17, 18, 16, 19, 20] as well as those based on non-volatile memory (NVM) devices such as Phase Change Memory (PCMs), Resistive RAMs (RRAMs), Spin-Transfer Torque RAMs (STT-RAMs), and Spin-Orbit Torque Magnetic RAMs (SOT-MRAMs) have been explored for implementing frequentist deep learning models [21, 14, 22]. Unlike SRAM IMC implementations, NVM devices allow multi-level conductance for multi-bit parameter storage, making them area and power-efficient.
However, NVM devices also pose several challenges for IMC MVM implementation. The high level of state-dependent programming [23] results in write error, and conductance drift results in instability of programmed weights [23, 10], severely impacting the classification accuracy. Additional sources of error include MVM errors due to IR drops on source-line due to high current in the all-parallel operation mode [18], and non-linear dependence of device current on applied voltage [24]. Furthermore, Analog-to-Digital converters (ADCs) are necessary to convert the analog output of the MVM operation to multi-bit digital values, and this impacts both the overall power consumption and area efficiency [18, 12].
Related works: We now discuss proposed IMC solutions in the literature that leverage the inherent stochasticity of NVM devices with the goal of implementing ensembling methods. An RRAM array-based Bayesian network with training implemented via Markov Chain Monte Carlo (MCMC) was demonstrated in [25], though the need for training in situ is time and resource intensive [26]. Bernoulli distributed noise of Domain Wall Magnetic Tunnel Junction (DW-MTJ) devices and SOT-MRAM were used in [21] and [9] respectively to generate Gaussian random variables. Since the weights are Gaussian distributed, two parameters – the mean and standard deviation – are required for each weight. In addition, generating a single Gaussian variable using the central limit theorem requires contributions from multiple noise sources, leading to expensive hardware overhead.
@@ I-C Main Contributions
  • We propose a crossbar array-based architecture that does not require complex ADCs. It employs a novel scheme that generates binary weights based on the simultaneous summation of conductances from the weight plane and noise plane. This permits synaptic weights to be generated on-the-fly in a power and area-efficient manner, enabling an ensembling-in-time implementation.
  • Unlike traditional IMC accelerators, our approach streams the inputs directly into the pre-neuron accumulator, avoiding expensive DACs in every row of the crossbar. With this approach, read pulses with fixed amplitude and pulse width can be applied on the bit-line for sampling, thus avoiding MVM errors due to the non-linear dependence of the device conductance on the applied voltage. This is markedly different from conventional IMC methods that require streaming the input bits of the activation variable bit-by-bit and performing shift-add operations to obtain MVM output [31, 10].
@@ II-A Binary Bayesian Neural Networks
Bayesian neural networks (BNNs) predict an output $\boldsymbol{y}$ for a given input $\boldsymbol{x}$ by evaluating the posterior distribution $p(\boldsymbol{w}|\mathcal{D})$, where the posterior distribution quantifies the likelihood of $\boldsymbol{w}$ under the evidence presented by the training dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i}$. At inference time, the contribution of each plausible model $p(\boldsymbol{y}|\boldsymbol{x},\boldsymbol{w})$ is weighted by the posterior $p(\boldsymbol{w}|\mathcal{D})$, yielding the marginal distribution [32, 29]
$p(y|\boldsymbol{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{w}|\mathcal{D})}\left[p(y|\boldsymbol{x},\boldsymbol{w})\right].$ (1)
In this work, we focus on the implementation of BBNNs, where each weight $w$ is restricted to a binary set of values $w\in\{-1,1\}$. Accordingly, the parameter space is defined as $\mathcal{W}=\{-1,1\}^{|\boldsymbol{w}|}$, where $|\boldsymbol{w}|$ represents the number of synaptic weights. Intuitively, BBNNs implement a set of neural networks where the weights are trained to follow a posterior distribution $p(\boldsymbol{w}|\mathcal{D})$ rather than a single point-estimate [29]. BBNNs are then trained through mean-field variational inference (VI) as discussed in [33, 34], where the posterior distribution $p(\boldsymbol{w}|\mathcal{D})$ is approximated by a parameterized distribution $q(\boldsymbol{w}|\boldsymbol{\lambda})$, where variational parameter $\boldsymbol{\lambda}$ consists of natural parameters of Bernoulli distribution $\lambda_{w}$ in the exponential family, i.e $\boldsymbol{\lambda}=\{\lambda_{w}\}_{w\in\boldsymbol{w}}\in\mathbb{R}^{|\boldsymbol{w}|}$. The probability parameter $p_{w}$ describing the distribution of binary weights $w$ is encoded as
@@
Training is performed using the BiNN optimizer, which is described in [34], with ELBO criterion (3) as the optimization problem to obtain a collection of synaptic probabilities $\boldsymbol{p}_{\boldsymbol{w}}=\{p_{w}\}_{w\in\boldsymbol{w}}$ from (2) (further details on training are provided in Supplementary Section A). Using an appropriate binary random variable generator, $\boldsymbol{p}_{\boldsymbol{w}}$ can be leveraged to create an ensemble of networks with $\boldsymbol{w}\sim q(\boldsymbol{w}|\boldsymbol{\lambda})$. As illustrated in Fig. 1, during inference, the outputs $p(y|\boldsymbol{x},\boldsymbol{w})$ of the sampled networks are aggregated to obtain the final marginal predictor
$p(y|\boldsymbol{x},\mathcal{D})\approx\frac{1}{N_{MC}}\sum_{i=1}^{N_{MC}}p(y|\boldsymbol{x},\boldsymbol{w}_{i}),$ (4)
@@ II-B Expected calibration error (ECE)
The ECE indicates the in-distribution calibration of a trained network by quantifying the expected discrepancy between accuracy and the confidence scores obtained by the output probability of the predicted class. For a given dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N_{s}}$ of $N_{s}$ samples, we denote as $B_{m}=\{(\hat{p}_{i},\hat{y}_{i},y_{i})\}_{i|\hat{p}_{i}\in I_{m}}$ the set of confidence-label pairs for which the predicted confidence $\hat{p}_{i}=p(y_{i}|x_{i},\mathcal{D})$ lies inside $I_{m}$, with $m\in\{1,...,N_{b}\}$. Here $N_{b}$ denotes the number of equally spaced intervals in $[0,1]$ and $I_{m}=\big{(}(m-1)/N_{b},m/N_{b}\big{]}$. The ECE is mathematically described, as per [4], as
@@ II-C Aleatoric and Epistemic Uncertainty
where $\{\boldsymbol{w}_{i}\}_{i=1}^{N_{MC}}\overset{\mathrm{i.i.d.}}{\sim}q(\boldsymbol{w}|\boldsymbol{\lambda})$ and $p(y|\boldsymbol{x},\boldsymbol{w}_{i})$ is the probability value for the given class $y\in\mathscr{C}$. This total uncertainty $U_{tot}$ can be decomposed into the uncertainty due to the inherently stochastic nature of the covariate-label pairs $(\boldsymbol{x},y)$, i.e., the aleatoric uncertainty; and the uncertainty stemming from the limited amount of evidence $\mathcal{D}$ i.e., the epistemic uncertainty.
@@
Ideally, for an out-of-distribution (OOD) data, the ensemble of predictors will disagree on predictions, causing a large epistemic uncertainty. In contrast, on in-distribution (IND) data, there should be a broad agreement on the prediction probabilities, and thus the epistemic uncertainty should be low. Fig. 2 illustrates the distinct roles that aleatoric and epistemic uncertainties play in understanding the predictions of a BBNN.
Fig. 2 and Fig. 2 consider the output of the BBNN to IND data. In Fig. 2, the models assign a relatively high probability score to one class (‘Truck’) and low scores for the rest, indicating low aleatoric uncertainty $(U_{a})$. In Fig. 2, the model assigns low and relatively similar probability scores to all classes, indicating higher $U_{a}$. The epistemic uncertainty ($U_{e}$) is low in both cases as there is agreement amongst the predictors on the outputs. This is indicated by error bars that signify the spread of probability scores per class across the ensemble. Fig. 2 and Fig. 2 consider OOD data corresponding to different types of inputs as compared to the trained data. In this case the high epistemic uncertainty is manifested by the disagreement amongst the probability distributions in the ensembles illustrated by a larger spread of probability values indicated by error bars due to the unfamiliar nature of the inputs.
@@ III Architecture and Design
In this section, we discuss the features of Bayes2IMC, our proposed architecture illustrated in Fig. 1. We first introduce the mathematical techniques that allow the utilization of device stochasticity as a source of noise distribution in Sec. III-A and the corresponding hardware implementation in Sec. III-B. We then describe the details of the core operations of the proposed Bayes2IMC architecture (Sec. III-C) and introduce the proposed output logit correction scheme (Sec. III-D). Finally, Sec. III-E discusses the drift compensation mechanism.